OpenStudy (anonymous):

how would you solve for x; 1/2 + 1/4 = 1/x

OpenStudy (anonymous):

add on the left then take the reciprocal

OpenStudy (anonymous):

Could you show me a example please?

OpenStudy (anonymous):

first solve 1/2+1/4

OpenStudy (anonymous):

aka add

OpenStudy (anonymous):

just add it do it in cals u'll find it easy

OpenStudy (anonymous):

3/4 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now you make x = to 3/4

OpenStudy (anonymous):

\[1/2+1/4=1/x\] \[0.75=1/x\] \[x=1/0.75\] \[x=1.333\] actually 0.75 is in decimal form of 3/4

OpenStudy (anonymous):

did u get it

OpenStudy (anonymous):

Yes thank you sheelo_mughal

OpenStudy (anonymous):

can i have a medal

OpenStudy (anonymous):

How do I give medals I'm new here

OpenStudy (anonymous):

just click best response on right side

OpenStudy (anonymous):

Oh thanks